EXERCISE 3. Suppose X;s are i.i.d. Beta B(3,6) (a) Compute E(5Xị - X1X2 – 4); Inn
Answer #1

page) @ we have Xis are iid Beta (316) we have to find EC5X, - xix 2-4) consider, pdof Beta BC.) fa) = ( -1 (1-2)- 0<x<! 7 BPage 2] using above formula iwe can arite, B(5,6) B(3.6) But B(4,) = 1B 18+ :. Ecxi) = 15.16 15+6 1346 131_ 15+ Ca (n-1):) 3.page 3 = tu.lg 13.10 - 31 x 8! 21xg! -0.3333 ) = 0.3333 TE(x Hly , find 6- f(x)= E(X2), 3-1 xoa- I (1- B (3,6) 2 OCX 221 othepage 4 Now, Consider E(5x; -*1x2-4), E (5xi-X1X2 - 4) - E(5 xi)- E (x,x)) - EC4) xis are independent, 1.e. Elx, x2) = E(X) •